OpenStudy (anonymous):

a contractor estimates that three person could repair house in 4 days if he uses 4 person instead of three how long should they take to complete the job?

OpenStudy (anonymous):

Let W be the amount of work to be done, and x be the rate of work for a person. Then we get: \[ \frac{W}{3x} = 4 \] Solve for x in terms of W: \[ x = \frac{W}{12} \] If he uses 3 people to do the same amount of work W, we can express the amount of time needed 'T' as: \[ T = \frac{W}{4x} \] Substituting the expression for x, you will get the answer T.